OpenStudy (anonymous):

Solve –3(2x – 1) = x + 10. You must show all work

OpenStudy (nubeer):

multiply -3 with the elecments inside the bracker -6x+3 = x+10 rest is simple algebra.. can u do from here?

OpenStudy (anonymous):

yes, I think so. Next you would add 3 to both sides... right?

OpenStudy (nubeer):

hmm add or subtract?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

That way it would cancel out -3

OpenStudy (nubeer):

-6x+3 = x+10 this is the equation.. now look again..

OpenStudy (anonymous):

oh! I wrote it down wrong... ok so you would subtract. It would then come out to be -6x=x+7 now would I multiply?

OpenStudy (nubeer):

hmm not multiply now u have to make all the x's on one side.

OpenStudy (anonymous):

so I would divide x to both sides? :/

OpenStudy (nubeer):